OpenStudy (anonymous):

Find the exact value of the trigonometric function. sin135

OpenStudy (anonymous):

\[\huge \sin135^0 = \sin (90^0 + 45^0) =\cos 45^0 =\frac{1}{\sqrt 2}\]

OpenStudy (kainui):

I prefer to visualize it as:  So you can think of it as just sin(45)